OpenStudy (anonymous):

An unknown compound contains only carbon, hydrogen, and oxygen (CxHyOz). Combustion of 2.50g of this compound produced 3.67g of carbon dioxide and 1.50g of water. How many moles of carbon, C, were in the original sample?

OpenStudy (anonymous):

@doc.brown Help please :D!

OpenStudy (anonymous):

>.>

OpenStudy (doc.brown):

The product weighs more than what you started with (3.67g+1.50g > 2.50g) so something is being added. Since we're talking about carbon diOXIDE, dihyrdogen OXIDE, and combustion, we can safely guess it's oxygen, so, \[CHO + O_{2} \rightarrow CO_{2} + H_{2}O\] You can see nothing else but our compound is contributing carbon to the carbon dioxide, so the number of carbon atoms there equals the number of carbon atoms in the original sample. Do you have your periodic table? Look at the amu for Carbon 12.011g/mol, and Oxygen 15.999g/mol. \[CO_{2}: (12.011\frac{g}{mol})\times2(15.999\frac{g}{mol})=44.009\frac{g}{mol}\]\[3.67g CO_{2} \times \frac{mol}{44.009g}=0.0834mol\] In 0.0834mol CO2 we have 2(0.0834 mol O) = 0.167mol O, and 0.0834mol C or if your teacher prefers scientific notation \[8.34\times10^{-2}molC\]